What is the solution to the problem 8-3(x-4)=4

Answer by trmbneprson(1)

8-3(x-4)=4
First, you distribute the -3 to (x-4): -3*x=-3x and -3*-4=12
Which equals: 8-3x+12=4
Second, you add like terms, 8 and 12: 8+12=20
Which equals: 20-3x=4
Then, you cancel out 20 to both sides: 20-20-3x=4-20
Which equals: -3x=-16
Then you cancel out the -3 by dividing: -3x/-3=-16/-3
Which equals: x=5.3 repeating